40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2624     Entry time: Mon Feb 22 11:38:05 2010
Author: joe, jenne and steve 
Type: Configuration 
Category: VAC 
Subject: vacuum is back to normal 

Morning condition: vacuum rack power is still off, no MEDM screen reading.....meaning unknown vacuum pressure.We closed PSL shutter immediately.

Joe restored c1iscepis and Jenne powered up the vac-rack UPS. Now the rest of the vac-rack power were restored from starting at the top to bottom.

P1 was reading 15 mTorr.  We restarted pumps and  set vacuum valve positions. V1 opening required Rob's recipe of elog # 1863 to defeat interlock that

has a non communicating gauge: PTP1

CC1 pressure just reached 1e-6 Torr at VAC NORMAL configuration.

ELOG V3.1.3-